OpenStudy (anonymous):

Use the Addition Formula for Sine to prove the Double-Angle Formula for Sine. sin 2x = 2 sin(x) cos(x) Rewrite 2x as x + x, and use the Addition Formula for Sine to simplify. sin 2x= sin(x + x) I have sin (x) cos(x) + cos (x) sin(x). I can't figure out what it EQUALS to.

OpenStudy (anonymous):

it equals 2sin(x)cos(x)

OpenStudy (anonymous):

\[\large \sin(x)\cos(x) + \cos(x)\sin(x) = 2\sin(x)\cos(x)\]

OpenStudy (anonymous):

Use an appropriate Half-Angle Formula to find the exact value of the expression. cos 112.5°